MICROWAVE AND MMWAVE STUDY OF CH_{3}SiH_{3} INCLUDING THE PERTURBATION-ALLOWED TORSION-VIBRATION DIFFERENCE BAND (v_{12} = 0, v_{6} = 3) ↔ (v_{12} = 1, v_{6} = 0)

\underline{I. OZIER}, Department of Physics and Astronomy, University of British Columbia, Vancouver, BC, Canada V6T 1Z1; and Laboratorium für Physikalische Chemie, Eidgenössische Technische Hochschule, CH-8093 Zürich, Switzerland; A. BAUDER, Laboratorium für Physikalische Chemie, Eidgenössische Technische Hochschule, CH-8093 Zürich, Switzerland; W. L. MEERTS, Department of Molecular and Laser Physics, NSRIM, University of Nijmegen, NL-6500 GL Nijmegen, The Netherlands; N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, Calgary, AB, Canada T2N 1N4; J. SCHRODERUS, Department of Physical Sciences, University of Oulu, FIN-90014, Linnanmaa, Finland.

The vibration-torsion-rotation Hamiltonian in CH_{3}SiH_{3} has been investigated using Fourier transform microwave methods and tunable sideband far-infrared spectroscopy. Four different studies have been carried out. First, the Q-branch of the torsion-vibration difference band (v_{12} = 0, v_{6} = 3) ↔ (v_{12} = 1, v_{6} = 0) has been measured between 17.8 and 26.6 GHz. When three quanta of the torsional mode ν_{6} are excited in the ground vibrational state (gs) for (σ = −1) torsional sublevels with K = 6, these transitions become allowed through resonant Coriolis-like coupling to the lowest lying degenerate mode ν_{12} with no quanta of ν_{6} excited. Second, direct l-doubling transitions in the state (v_{12} = 1, v_{6} = 0) have been observed between 8.3 and 17.5 GHz for both σ = 0 and σ = ±1. In the limit that the intervibrational interactions vanish, the σ-splitting between lines of the same J would be difficult to resolve, but frequency differences of more than 1 GHz due to these interactions have been determined. Third, the (J = 1 ← 0) spectrum just below 22 GHz has been re-measured with higher resolution for 0 ≤ v_{6} ≤ 4 in the gs and for (v_{6} = 0) in ν_{12}. Finally, the (J = 45 ← 44) spectrum near 1 THz has been obtained for 0 ≤ v_{6} ≤ 2 in the gs. A global data set of 3423 frequencies has been formed by merging the present 123 measurements with the data set used recently in the simultaneous analysis of the ν_{12} and ν_{5} bands. By refining the (gs/ν_{12}/ν_{5}) Hamiltonian developed in this earlier work, a good fit to within experimental error has been obtained by varying 45 parameters. The grouping of the torsional motion with rotational rather than vibrational degrees of freedom will be discussed.